A balloon initially contains 7 gram of Nitrogen and then 14 gram of nitrogen is added to the balloon expand its volume to 12 litre at the same temperature and pressure. Find the initial volume of the balloon.

Grade:10

1 Answers

Arun
6 years ago
Dear student
 
For an ideal gas
PV = nRT
P* V1 = n1 *R*T
 
P* V1 = (w1/m1) RT
P V1 = (7/28)*RT         .........(1)
 
Now
P*12 = (14/28) *RT                 .....(2)
 
Now
On dividing equation 1 by equation 2.
V1 = 6 litres
 
Hence initial volume = 6 litres.
 
 
Regards
Arun (askIITians forum expert)